How much pure acid should be mixed with 6 gallons of 50% acid solution in order to get an 80% acid solution?
Answer by richwmiller(17219)

We have 6 gal of 50% acid
We want to add some 100% acid to make 80% acid
1*x+0.5*6=0.8(6+x)
1*x+3=4.8+0.8x
1x-0.8x=4.8-3
0.2x=1.8
x=9 gal at 100% (9) is added to 6 gal at 50%(3) to make 80% of 15 gal(12)
check
1*9+0.5*6=0.8(6+9)
9+3=0.8(15)
12=12
ok